# Test-Time Adaptation of Vision-Language Models for Open-Vocabulary Semantic Segmentation

### v1

@@ 5.2 Final Comparison with Alternative Adaptation Methods
Performance on Clean Data (No Distributional Shift).  We begin by evaluating MLMP on clean test data (original), where no distributional shift is present. This setting is crucial, as TTA methods must avoid degrading performance when adaptation is unnecessary. As shown in Table 3, MLMP achieves strong mIoU gains of +7.85, +5.66, +3.72, and +3.04 on V20, V21, P59, and P60, and +5.04/+2.91 on challenging datasets COCOObject and COCOStuff. In contrast, most alternative adaptation methods fail to improve performance in this setting. These gains highlight the robustness and generalization of our method, even when no explicit domain shift is present.
Performance Under Distributional Shift.  Under distributional shifts, the advantages of MLMP become even more apparent. As shown in Table 3, MLMP consistently outperforms both the zero-shot baseline and existing adaptation methods, achieving mIoU gains of +8.60, +5.50, +3.15, and +2.68 on V20-C, V21-C, P59-C, and P60-C, respectively. Beyond standard corruptions, we further evaluate on the Cityscapes dataset, which presents natural domain shifts such as environmental variation, weather conditions, and resolution differences. Despite its challenging nature and low zero-shot performance, MLMP improves mIoU by +3.86, demonstrating its real-world adaptability. To push this further, we apply 15 corruption types to create Cityscapes-C, where MLMP still yields a +1.39 gain. While TENT provides modest improvements, most other adaptation methods—including ClipArTT, TPT, and WATT—either fail to improve or degrade performance. These results highlight that naive strategies like pseudo-labeling, prompt tuning, or weight averaging are insufficient for open-vocabulary segmentation, and emphasize the need for segmentation-specific adaptation techniques such as MLMP. Detailed results can be found in Appendix H.
@@ 6 Conclusion
We presented MLMP, a plug-and-play test-time adaptation framework for open-vocabulary semantic segmentation that can be integrated with any OVSS method. By combining uncertainty-aware fusion of intermediate ViT features with a novel loss-level integration of multiple prompt templates, MLMP consistently enhances performance across both clean and shifted domains—including common corruptions and natural distributional shifts. Our comprehensive OVSS-TTA benchmark—covering seven datasets and 82 distinct test scenarios—demonstrates MLMP’s broad applicability and establishes a rigorous evaluation protocol for future work in adaptive, language-aware segmentation. While MLMP demonstrates strong, consistent gains, there remain opportunities to further refine its components. In particular, our current layer-weighting mechanism relies on entropy estimates from a shared projection head, which may not fully reflect each layer’s unique characteristics. Future work could investigate more flexible architectures—such as lightweight adapters or dedicated projection modules per layer—to more accurately assess and fuse intermediate features.
@@ Appendix B Implementation Details
Unless otherwise noted, all experiments utilize NACLIP [30] with ViT-L/14 as the OVSS backbone. We adapt only the LayerNorm parameters within the vision encoder, amounting to approximately 0.02% of the model’s total parameters. Our adaptation setup follows prior work in classification [5, 31], using the Adam optimizer with a fixed learning rate of $0.001$ and 10 adaptation steps (iterations) across all experiments. Additionally we use batch of 2 images during adaptation. After each batch, model weights are explicitly reset to their original pre-adaptation values to ensure that each batch is adapted independently, without leveraging information from previously processed data. Following standard settings from  [1], we use the default softmax temperature value of $100$ in all experiments. All images are resized to $224\times 224$ pixels. Due to the high resolution of images in the Cityscapes dataset, we split them into overlapping patches of size $224\times 224$ pixels with an overlap of 112 pixels between patches. The segmentation predictions from these patches are aggregated to reconstruct the final, full-resolution segmentation maps. No image augmentation techniques are applied during either the adaptation or evaluation phases.
@@
We adapt other baseline methods (TENT [8], TPT [6], WATT [5], CLIPArTT [31]) to work with spatial tokens. General adaptation hyperparameters (optimizer, learning rate, adaptation steps, and batch size) remain consistent with our setup. Method-specific hyperparameters or components are retained as reported in their original implementations. Specifically, for WATT, we used the sequential version (WATT-S) with default values of $l=2$ and $m=5$, for CLIPArTT we used the default $k=3$, and for TPT we used the 4 learnable tokens. Adapting these baseline methods to the segmentation task allowed us to systematically evaluate how various adaptation strategies—such as prompt tuning (TPT), pseudo-labeling (WATT, CLIPArTT), prompt refinement (CLIPArTT), and weight averaging (WATT)—perform in the context of open-vocabulary segmentation.
@@ Appendix C Template Details
More specifically, for the ablation in Fig. 4 of the main paper, we vary the number of templates $T$ as follows:
@@ Appendix D Dataset Details
This section details the datasets used in our experiments, along with the synthetic corruptions applied to evaluate robustness under distribution shifts.
@@
  • Pascal VOC (V20/V21) [34]: Contains 20 foreground object classes (v20) with a background class (v21), widely used for benchmarking semantic segmentation tasks.
  • Pascal Context (P59/P60) [35]: An extension of the Pascal VOC 2010 dataset, providing pixel-level annotations for more than 400 classes. Due to sparsity, a frequently used subset includes 59 object classes plus a background class, totaling 60 categories.
  • Cityscapes [36]: A large-scale dataset for semantic segmentation of urban street scenes. It comprises around 5,000 finely annotated images from 50 cities, recorded under various daylight conditions, featuring dynamic objects, varying layouts, and changing backgrounds, capturing significant natural domain shifts.
  • COCO-Stuff [37]: Extends the original COCO dataset by adding annotations for background categories ("stuff"), resulting in 171 classes—80 objects, and 91 stuff categories.
  • COCO-Object [38]: A subset of the original COCO dataset, consisting exclusively of 80 object categories without background annotations.
In addition to the original versions of each dataset—some of which already reflect natural domain shifts (e.g., Cityscapes)—we further assess the robustness of all methods by applying synthetic corruptions. Inspired by ImageNet-C [33], we apply 15 synthetic corruptions to evaluate the robustness of segmentation models under various perturbations. These corruptions include Gaussian noise, shot noise, impulse noise, defocus blur, glass blur, motion blur, zoom blur, snow, frost, fog, brightness variations, contrast variations, elastic transformations, pixelation, and JPEG compression. Each corruption is applied at severity level 5, representing the most challenging scenario.
We resize images from all datasets to $224\times 224$ pixels. Due to the high resolution of Cityscapes images, we process them as overlapping patches of size $224\times 224$ pixels (with overlaps of 112 pixels). Predictions for these patches are subsequently aggregated to produce the final segmentation maps.
This curated set of 82 distinct test scenarios rigorously benchmarks our method across a comprehensive spectrum of conditions. It encompasses not only typical natural domain shifts and synthetic corruptions but also significant variations in image resolution, scene complexity, object scale, and semantic granularity.
@@ Appendix E Computational Complexity
In terms of latency, MLMP completes both adaptation and evaluation in just 0.582 ms, which is only marginally slower than TENT (0.480 ms) but substantially faster than WATT (5.215 ms) and CLIPArTT (3.525 ms). Despite leveraging multi-level fusion and multiple prompt templates, MLMP maintains a lightweight computational profile with only 82.4 GFLOPs and 82.9 GFLOPs for adaptation and evaluation, respectively—comparable to TENT and significantly lower than all other baselines. Furthermore, MLMP’s peak memory usage is among the lowest at 2,093.9 MB, and like TENT, it updates only 0.02% of the model parameters. These results highlight MLMP’s efficiency: it delivers rich representational capacity through multi-level and multi-prompt integration by boosting the results significantly (as shown in Table 4 of the main paper), yet remains almost as lightweight as the simplest baseline.
Table 5: Computational-complexity comparison across methods. For GFLOPs, only the forward-pass cost in adaptation and evaluation is measured; by common practice, the cost of back-propagation in adaptation phase can be approximated as twice the forward cost. A ✓ in the second column indicates that all text information are encoded once and cached. A dagger ($\dagger$) indicates that TPT adds additional parameters beyond the original network.
Method One-time Text Encoder Time (sec.) $\downarrow$ GFLOPs $\downarrow$ Max Memory (MB) $\downarrow$ Learn. Params (Ratio) $\downarrow$
@@ G.1 Comparison Across ViT Backbones
To evaluate whether our method generalizes across different vision transformer backbones, we replicate the main experiments using ViT-B/16 and ViT-B/32 in place of the default ViT-L/14 model. These backbones represent lighter configurations, with fewer parameters and larger patch (32 and 16). As shown in Tables 7 and 8, MLMP continues to provide substantial improvements over all baselines across both configurations, on both clean data (V20 Original) and under severe synthetic corruptions. These results confirm that the benefits of our multi-level and multi-prompt adaptation strategy are not tied to model scale or specific architectural configurations, and remain effective even in lower-resolution settings.
@@ G.2 Comparison Across OVSS Methods
Specifically, we consider the original CLIP [1], adapted for pixel-wise segmentation via patch-level similarity, and SCLIP [27], which incorporates spatial priors into the vision-language matching process. As shown in Table 9, applying MLMP on top of these OVSS baselines yields consistent improvements across both clean and corrupted settings.
Table 9: Performance comparison of test-time adaptation methods using the original CLIP [1] and SCLIP [27] as the OVSS model with a ViT-B/16 backbone. MLMP consistently improves performance across all settings, highlighting its strong generalization capabilities.
OVSS: CLIP [1] Adaptation Method
@@
V21-C Average 28.02 45.33 54.80
OVSS: SCLIP [27] Adaptation Method
@@ G.3 Generalization to Emerging VLMs
To further assess the generality of our approach, we evaluate MLMP on SigLIP v2 [39], one of the most recently introduced vision–language models. As shown in Table 10, MLMP continues to deliver consistent improvements across both natural and corrupted datasets, indicating that the core components of our method—multi-level and multi-prompt aggregation—generalize well beyond CLIP-based architectures.
@@
V20-C Average 40.93 42.33
Appendix H Visualization of Layer-Wise Confidence Weights

### v2

@@ 5.2 Final Comparison with Alternative Adaptation Methods
Performance on Clean Data (No Distributional Shift).  We begin by evaluating MLMP on clean test data (original), where no distributional shift is present. This setting is crucial, as TTA methods must avoid degrading performance when adaptation is unnecessary. As shown in Table 5.1, MLMP achieves strong mIoU gains of +7.85, +5.66, +3.72, and +3.04 on V20, V21, P59, and P60, and +5.04/+2.91 on challenging datasets COCOObject and COCOStuff. In contrast, most alternative adaptation methods fail to improve performance in this setting. These gains highlight the robustness and generalization of our method, even when no explicit domain shift is present.
Performance Under Distributional Shift.  Under distributional shifts, the advantages of MLMP become even more apparent. As shown in Table 5.1, MLMP consistently outperforms both the zero-shot baseline and existing adaptation methods, achieving mIoU gains of +8.60, +5.50, +3.15, and +2.68 on V20-C, V21-C, P59-C, and P60-C, respectively. Beyond standard corruptions, we further evaluate on the Cityscapes dataset, which presents natural domain shifts such as environmental variation, weather conditions, and resolution differences. Despite its challenging nature and low zero-shot performance, MLMP improves mIoU by +3.86, demonstrating its real-world adaptability. To push this further, we apply 15 corruption types to create Cityscapes-C, where MLMP still yields a +1.39 gain. While TENT provides modest improvements, most other adaptation methods—including ClipArTT, TPT, and WATT—either fail to improve or degrade performance. These results highlight that naive strategies like pseudo-labeling, prompt tuning, or weight averaging are insufficient for open-vocabulary segmentation, and emphasize the need for segmentation-specific adaptation techniques such as MLMP. Detailed results can be found in the Appendix.
Table 5: mIoU comparison on realistic (ACDC) and rendered (GTA-V) domain shifts. Full ACDC results (including reference/clean views) are provided in the Appendix.
OVSS: NACLIP Adaptation Method
Dataset No Adapt. TENT MLMP
ACDC Fog 23.88 26.89 $\pm$0.04 33.33 $\pm$0.04
Night 22.12 24.17 $\pm$0.00 24.76 $\pm$0.03
Rain 23.86 26.84 $\pm$0.04 32.44 $\pm$0.04
Snow 23.54 27.25 $\pm$0.05 30.59 $\pm$0.03
[…]
While Cityscapes already embodies natural distributional shifts caused by variations in lighting, camera viewpoint, and urban layouts, we further assess MLMP on datasets explicitly designed to capture targeted domain shifts. Specifically, we evaluate on ACDC acdc, which contains real-world adverse conditions (Fog, Night, Rain, and Snow), and GTA-V gtav, a photorealistic, game-rendered dataset that introduces a distinct synthetic distribution shift relative to real imagery seen during CLIP pre-training. As summarized in Table 5.2, MLMP achieves consistent improvements over both the non-adapted baseline and TENT across all ACDC domains, yielding on average +6 mIoU gains under real-world conditions. Similarly, on the GTA-V dataset, MLMP improves by +3.8 mIoU over the baseline, further confirming its robustness across both realistic and rendered distribution shifts.
@@ 6 Conclusion
We presented MLMP, a plug-and-play test-time adaptation framework for open-vocabulary semantic segmentation that can be integrated with any OVSS method. By combining uncertainty-aware fusion of intermediate ViT features with a novel loss-level integration of multiple prompt templates, MLMP consistently enhances performance across both clean and shifted domains—including common corruptions and natural distributional shifts. Our comprehensive OVSS-TTA benchmark—covering nine datasets and 87 distinct test scenarios—demonstrates MLMP’s broad applicability and establishes a rigorous evaluation protocol for future work in adaptive, language-aware segmentation. While MLMP demonstrates strong, consistent gains, there remain opportunities to further refine its components. In particular, our current layer-weighting mechanism relies on entropy estimates from a shared projection head, which may not fully reflect each layer’s unique characteristics. Future work could investigate more flexible architectures—such as lightweight adapters or dedicated projection modules per layer—to more accurately assess and fuse intermediate features.
Test-Time Adaptation of Vision-Language Models for Open-Vocabulary Semantic Segmentation - Appendix
@@ Appendix B Implementation Details
Unless otherwise noted, all experiments utilize NACLIP [naclip] with ViT-L/14 as the OVSS backbone. We adapt only the LayerNorm parameters within the vision encoder, amounting to approximately 0.02% of the model’s total parameters. Our adaptation setup follows prior work in classification [watt, clipartt], using the Adam optimizer with a fixed learning rate of $0.001$ and 10 adaptation steps (iterations) across all experiments. Additionally we use batch of 2 images during adaptation. After each batch, model weights are explicitly reset to their original pre-adaptation values to ensure that each batch is adapted independently, without leveraging information from previously processed data. Following standard settings from  [clip], we use the default softmax temperature value of $100$ in all experiments. All images are resized to $224\times 224$ pixels. Due to the high resolution of images in the Cityscapes dataset, we split them into overlapping patches of size $224\times 224$ pixels with an overlap of 112 pixels between patches. The segmentation predictions from these patches are aggregated to reconstruct the final, full-resolution segmentation maps. No image augmentation techniques are applied during either the adaptation or evaluation phases.
@@
We adapt other baseline methods (TENT [tent], TPT [tpt], WATT [watt], CLIPArTT [clipartt]) to work with spatial tokens. General adaptation hyperparameters (optimizer, learning rate, adaptation steps, and batch size) remain consistent with our setup. Method-specific hyperparameters or components are retained as reported in their original implementations. Specifically, for WATT, we used the sequential version (WATT-S) with default values of $l=2$ and $m=5$, for CLIPArTT we used the default $k=3$, and for TPT we used the 4 learnable tokens. Adapting these baseline methods to the segmentation task allowed us to systematically evaluate how various adaptation strategies—such as prompt tuning (TPT), pseudo-labeling (WATT, CLIPArTT), prompt refinement (CLIPArTT), and weight averaging (WATT)—perform in the context of open-vocabulary segmentation.
@@ Appendix C Template Details
More specifically, for the ablation in Figure 5 of the main paper, we vary the number of templates $T$ as follows:
@@ Appendix D Dataset Details
This section details the datasets used in our experiments, along with the synthetic, real, and rendered domain shifts applied to evaluate robustness under distributional changes.
@@
  • Pascal VOC (V20/V21) [PascalVoc]: Contains 20 foreground object classes (v20) with a background class (v21), widely used for benchmarking semantic segmentation tasks.
  • Pascal Context (P59/P60) [PascalContext]: An extension of the Pascal VOC 2010 dataset, providing pixel-level annotations for more than 400 classes. Due to sparsity, a frequently used subset includes 59 object classes plus a background class, totaling 60 categories.
  • Cityscapes [Cityscapes]: A large-scale dataset for semantic segmentation of urban street scenes. It comprises around 5,000 finely annotated images from 50 cities, recorded under various daylight conditions, featuring dynamic objects, varying layouts, and changing backgrounds, capturing significant natural domain shifts.
  • COCO-Stuff [CocoStuff]: Extends the original COCO dataset by adding annotations for background categories ("stuff"), resulting in 171 classes—80 objects, and 91 stuff categories.
  • COCO-Object [CocoObject]: A subset of the original COCO dataset, consisting exclusively of 80 object categories without background annotations.
  • ACDC [acdc]: Includes pixel-level annotations for 19 semantic classes, covering diverse driving scenes under fog, night, rain and snow conditions.
  • GTA-V [gtav]: Contains  24,966 synthetic images with pixel-accurate semantic labels; one version aligns to 34/19 classes compatible with real-world segmentation benchmarks.
In addition to the original versions of each dataset—some of which already reflect natural domain shifts (e.g., Cityscapes)—we further assess the robustness of all methods by applying synthetic corruptions. Inspired by ImageNet-C [cifar10c&10.1], we apply 15 synthetic corruptions to evaluate the robustness of segmentation models under various perturbations. These corruptions include Gaussian noise, shot noise, impulse noise, defocus blur, glass blur, motion blur, zoom blur, snow, frost, fog, brightness variations, contrast variations, elastic transformations, pixelation, and JPEG compression. Each corruption is applied at severity level 5, representing the most challenging scenario.
We resize images from all datasets to $224\times 224$ pixels. Due to the high resolution of Cityscapes and ACDC images, we process them as overlapping patches of size $224\times 224$ pixels (with overlaps of 112 pixels). Predictions for these patches are subsequently aggregated to produce the final segmentation maps.
This extended benchmark comprises a total of 87 distinct test scenarios, encompassing synthetic, real, and rendered domain shifts. It provides a rigorous and comprehensive evaluation protocol that captures variations in resolution, scene diversity, object size, and semantic granularity.
@@ Appendix E Computational Complexity
In terms of latency, MLMP completes both adaptation and evaluation in just 0.582 ms, which is only marginally slower than TENT (0.480 ms) but substantially faster than WATT (5.215 ms) and CLIPArTT (3.525 ms). Despite leveraging multi-level fusion and multiple prompt templates, MLMP maintains a lightweight computational profile with only 82.4 GFLOPs and 82.9 GFLOPs for adaptation and evaluation, respectively—comparable to TENT and significantly lower than all other baselines. Furthermore, MLMP’s peak memory usage is among the lowest at 2,093.9 MB, and like TENT, it updates only 0.02% of the model parameters. These results highlight MLMP’s efficiency: it delivers rich representational capacity through multi-level and multi-prompt integration by boosting the results significantly (as shown in Table 5.1 of the main paper), yet remains almost as lightweight as the simplest baseline.
Table 7: Computational‐complexity comparison across methods. For GFLOPs, only the forward‐pass cost in adaptation and evaluation is measured; by common practice, the cost of back‐propagation in adaptation phase can be approximated as twice the forward cost. A ✓ in the second column indicates that all text information are encoded once and cached. A dagger ($\dagger$) indicates that TPT adds additional parameters beyond the original network.
Method One‐time Text Encoder Time (sec.) $\downarrow$ GFLOPs $\downarrow$ Max Memory (MB) $\downarrow$ Learn. Params (Ratio) $\downarrow$
@@ G.1 Comparison Across ViT Backbones
To evaluate whether our method generalizes across different vision transformer backbones, we replicate the main experiments using ViT-B/16 and ViT-B/32 in place of the default ViT-L/14 model. These backbones represent lighter configurations, with fewer parameters and larger patch (32 and 16). As shown in Tables 9 and 10, MLMP continues to provide substantial improvements over all baselines across both configurations, on both clean data (V20 Original) and under severe synthetic corruptions. These results confirm that the benefits of our multi-level and multi-prompt adaptation strategy are not tied to model scale or specific architectural configurations, and remain effective even in lower-resolution settings.
@@ G.2 Comparison Across OVSS Methods
Specifically, we consider the original CLIP [clip], adapted for pixel-wise segmentation via patch-level similarity, and SCLIP [wang2025sclip], which incorporates spatial priors into the vision-language matching process. As shown in Table 11, applying MLMP on top of these OVSS baselines yields consistent improvements across both clean and corrupted settings.
Table 11: Performance comparison of test-time adaptation methods using the original CLIP [clip] and SCLIP [wang2025sclip] as the OVSS model with a ViT-B/16 backbone. MLMP consistently improves performance across all settings, highlighting its strong generalization capabilities.
OVSS: CLIP [clip] Adaptation Method
@@
[…]
OVSS: SCLIP [wang2025sclip] Adaptation Method
@@ G.3 Generalization to Emerging VLMs
To further assess the generality of our approach, we evaluate MLMP on SigLIP v2 [siglip2], one of the most recently introduced vision–language models. As shown in Table 12, MLMP continues to deliver consistent improvements across both natural and corrupted datasets, indicating that the core components of our method—multi-level and multi-prompt aggregation—generalize well beyond CLIP-based architectures.
@@
[…]
Appendix H Effect of Longer Prompts
Recent studies such as Long-CLIP [longclip] and TULIP [tulip] have explored extending the text length capability of vision–language models, suggesting that richer linguistic descriptions may improve alignment. Motivated by this, we investigated whether incorporating longer and more descriptive prompts could further enhance MLMP’s performance.
To this end, we generated extended templates derived from our original seven class-agnostic prompts using ChatGPT, together with extended class names where each category was expressed in full natural language. For example, “a photo of the large {class}” becomes “a photograph of a very large {class}, where the size of the object dominates the frame or is shown in contrast to smaller elements,” while “aeroplane” becomes “a powered flying vehicle with fixed wings and engines, designed to transport people or cargo through the air over long distances.”
As summarized in Table 13, all experiments were conducted using the Long-CLIP [longclip] backbone. We evaluate three text variants: (1) the default short templates used throughout our main experiments, (2) Extended Templates, where each template is replaced with a longer and more descriptive sentence, and (3) Extended Classes, where category names are expanded into full natural-language descriptions. For each variant, we report results with and without our MLMP adaptation (+ MLMP). All scores correspond to mIoU.
Table 13: Effect of longer prompts and extended class descriptions on segmentation performance (mIoU). All experiments use Long-CLIP as the OVSS model. Each pair of columns shows results without and with MLMP adaptation.
OVSS: Long-CLIP Adaptation Setting
Dataset Default Templates Extended Templates Extended Classes
No Adapt. + MLMP No Adapt. + MLMP No Adapt. + MLMP
V20 39.64 54.45  $\pm$0.09 35.55 53.43  $\pm$0.06 25.66 43.92  $\pm$0.04
[…]
V21 16.83 19.54  $\pm$0.02 15.87 19.15  $\pm$0.04 11.27 15.25  $\pm$0.03
[…]
Overall, these results indicate that MLMP’s adaptation mechanism effectively handles both concise and extended textual inputs without overfitting to prompt verbosity. We include this analysis to provide a clearer picture of MLMP’s behavior under longer or more descriptive prompts.
Appendix I Effect of Adaptation Iterations
The number of adaptation iterations in TTA varies across the literature. While some studies evaluate performance under a single adaptation step, others perform multiple iterations to improve convergence. In our main experiments, we followed the common 10-iteration setup adopted in prior TTA works [tent, watt, clipartt], as it provides a consistent protocol for comparing vision–language adaptation methods such as CLIPArTT and WATT. For fairness, we used identical hyperparameters across all methods.
To further examine the influence of the iteration count, we also evaluated MLMP under a stricter one-iteration protocol, where only a single adaptation step is performed. As summarized in Table 14, MLMP still improves over the baseline and outperforms TENT even with just one iteration, which demonstrates the effectiveness of our method even under a stricter setting.
Table 14: Evaluation of MLMP under a single-iteration TTA protocol compared to standard baselines (mIoU). All experiments use NACLIP as the OVSS model.
OVSS: NACLIP Adaptation Method
Dataset No Adapt. TENT MLMP
V20 75.91 76.80  $\pm$0.01 79.88  $\pm$0.02
[…]
V21 45.12 45.36  $\pm$0.00 50.24  $\pm$0.01
[…]
Appendix J Episodic vs. Online Adaptation
We further investigate the difference between episodic (reset-based) and online adaptation settings in test-time adaptation (TTA). In our main experiments, all methods—including MLMP and baselines—were evaluated under the episodic setup, where the model is reset to its initial weights after each batch. This protocol avoids cumulative error propagation and is widely used in prior TTA works [tent, clipartt, watt].
To analyze the impact of continuous updates, we also tested an online version of MLMP, where model parameters are updated sequentially without resets and carried over to the next batch. We used identical hyperparameters as in the episodic setting, and further experimented with lower learning rates and fewer adaptation steps for stability.
As shown in Table 15, we observe that while the online variant improves over baselines, it still underperforms compared to the reset-based (episodic) setting. We believe this is primarily due to error accumulation and model drift, as documented in prior work showing that some recent classification TTA methods exhibit performance degradation in an online setting—even when the distribution is static [pitfalls]. This issue is particularly pronounced in semantic segmentation, where dense, spatial predictions make the model more sensitive to incorrect updates, and entropy minimization can amplify early misclassifications. Over time, this leads the model to drift away from its well-initialized source parameters, reducing its ability to generalize across the target domain.
Episodic (reset) adaptation avoids the runaway effects of carrying over errors, which is why it tends to be safer [2]. In contrast, fully online adaptation—though attractive for its potential to continuously refine the model—must confront these challenges.
We believe that online adaptation in segmentation can be improved through several future directions: (1) introducing occasional resets or using exponential moving average (EMA) of model weights to limit drift and reduce accumulated errors; (2) filtering high-entropy samples or gradients, which is particularly important in segmentation due to its spatial granularity and sensitivity to local noise; and (3) incorporating an auxiliary self-supervised objective, such as rotation prediction [1], to provide a more reliable adaptation signal—though this may come with increased computational cost. Notably, the fact that our method works effectively with very few test samples and does not rely on accumulating state makes it efficient for real-world applications.
Table 15: Comparison of episodic (reset-based) and online adaptation settings for MLMP and TENT (mIoU). All experiments use NACLIP as the OVSS model.
OVSS: NACLIP Adaptation Setting
Dataset No Adapt. Episodic: TENT Episodic: MLMP Online: TENT Online: MLMP
V20 75.91 77.00  $\pm$0.04 83.76  $\pm$0.00 76.44  $\pm$0.21 79.19  $\pm$0.45
[…]
Appendix K Visualization of Layer-Wise Confidence Weights